# Rational Linkages: From Poses to 3D-printed Prototypes

### v1

@@ 1 Introduction
Single-loop N-bar linkages are closed kinematic chains that connect a number of rigid bars via revolute or prismatic joints in one loop. Over-constrained mechanisms are a special class of such linkages. In this class, the Grübler-Kutzbach-Chebyshev formula, computing the degrees-of-freedom (DoF) of the mechanism, fails and the causes of this failure are special geometric properties of the design parameters of the mechanism. Because of the low number of DoFs, single-loop linkages can be synthesized as simple single-purpose devices for specific tasks, such as pick-and-place operations. This can be advantageous compared to serial robots, because they can perform complicated motions following space-curves of higher degree with a low number of active joints. Up to now, a major problem for the industrial application of these mechanisms is self-intersections [1], as discussed later.
Custom mechanisms designed for specific tasks are a wide research topic, with many studies conducted in fields such as robot manipulation [2, 3, 4] or locomotion [5, 6], whereby the design parameters are mostly obtained by numerical optimization. The tools presented in this paper, however, rely on algebraic methods. The problem of designing a spatial linkage analytically can be divided into the following steps: Pose interpolation, synthesis of the kinematic structure, and design of its physical realization, which also involves a collision analysis and CAD modeling to prepare 3D printing.
@@
The prime tasks in designing a custom mechanism or manipulator is the determination of the poses, which should be achieved by the end-effector tool of the device. The presented package implements for this purpose a method introduced by Hegedüs et al. [7], which interpolates four poses using a rational curve in the special Euclidean displacement group SE(3). When a rational motion curve on SE(3) is obtained, rational motion factorization introduced by Hegedüs, Schicho, and Schröcker [8, 9] is applied. The factors of the rational motion curve relate to 1-DoF revolute or prismatic joints, which can be connected to form a linkage capable of performing the synthesized end-effector motion.
@@ 2 Mathematical Background
A possibility to describe lines in 3D space using coordinates are the so called Plücker coordinates. These are six-dimensional vectors $\mathbf{l}=(g_{0},g_{1},g_{2},g_{3},g_{4},g_{5})$. The line’s direction is $\mathbf{g}=(g_{0},g_{1},g_{2})$, its moment vector $\overline{\mathbf{g}}=(g_{3},g_{4},g_{5})$ is obtained as $\overline{\mathbf{g}}=\mathbf{q}\times\mathbf{g}$ where $\mathbf{q}$ is an arbitrary point on the line. Plücker coordinates fulfill the Plücker condition $\mathbf{g}^{T}\cdot\overline{\mathbf{g}}=0$. More on Plücker coordinates can be found in Pottmann and Wallner [10, Chapter 2].
Study parameters or dual quaternions are used for the description of rigid body transformations either as 8-tuples $\mathbf{p}=(p_{0},p_{1},\ldots,p_{7})$ or by using two quaternions $\mathbf{a},\mathbf{b}$ as dual quaternion $\mathbf{p}=\mathbf{a}+\varepsilon\mathbf{b}$, where $\mathbf{a}=(p_{0},p_{1},p_{2},p_{3})$, $\mathbf{b}=(p_{4},p_{5},p_{6},p_{7})$ and $\varepsilon$ is the dual unit $\varepsilon^{2}=0$ (see Bottema and Roth [11, Chapter XIII]). Study parameters have to fulfill the quadratic equation $\mathbf{a}^{T}\cdot\mathbf{b}=0$, which is the equation of the so-called Study quadric. Points $\mathbf{q}$ and lines $\mathbf{l}$ can be embedded into the algebra of dual quaternions by $\mathbf{q}=(1,0,0,0,0,q_{x},q_{y},q_{z})$, where $(q_{x},q_{y},q_{z})$ are the coordinates of the point in the Euclidean space $E^{3}$ and $\mathbf{l}=(0,g_{0},g_{1},g_{2},0,g_{3},g_{4},g_{5})$ where $g_{i}$ are the Plücker coordinates of the line. Transformation $\mathbf{p}$ transforms a point or a line by
@@
Given a rigid body motion represented by the polynomial $C(t)$ over the dual quaternions, it is generically possible to synthesize a linkage that performs this motion by combining different factorizations of $C(t)$, c.f. Hegedüs, Schicho, and Schröcker [9].
We compute the factorizations of the motion polynomial with Python using the open source package biquaternion-py by Thimm [12]. This package provides a user-friendly implementation of a general dual quaternion algebra and polynomials over it for numerical and symbolical computations. It is ready for use, even by non-specialists but some attention needs to be paid to details of symbolic computations (providing extension fields over which polynomials factor if they cannot be inferred implicitly) and numerics (for example taking into account rounding errors in polynomial division). More about the general usage of this package and its intricacies can be found in the documentation of the biquaternion-py package [12].
@@ 3 Rational Mechanism Toolbox
While the previous section introduced methods and tools that are implemented but are already known, this section presents the main contribution of this work. It is the added functionality in terms of rapid prototyping of the linkages. Installation instructions can be found in the documentation hosted on the ReadTheDocs platform and are accessible at [13]. The source is available from Gitlab repository hosted by the University of Innsbruck [14], which also serves as the code maintenance server.
As an example to demonstrate the features, we will use a general Bennett mechanism that was synthesized by Brunnthaler et al. [15]. The motion curve of degree two of the given Bennett linkage has the equation specified by Study parameters
@@
c = RationalCurve(C(t)) normal-$\triangleright$ create an object Rational Curve from the input equations
factors = FactorizationProvider.factorize_motion_curve(c) normal-$\triangleright$ factorize curve
m = RationalMechanism(factors) normal-$\triangleright$ create a mechanism object
plt = Plotter.plot$(m)$ normal-$\triangleright$ plot the mechanism
@@
In the toolbox, a collision check for the line model is implemented. It uses dual quaternion operations on lines to create parametric equations of moving lines, joints and links, during a full cycle of the motion. The lines move relatively between each other and two lines $\mathbf{l_{0}}$ and $\mathbf{l_{1}}$ represented by their Plücker coordinates intersect if the following equation is fulfilled [10, p.140]:
@@
In the package, the command RationalMechanism.get_design(scale) generates design parameters for a pre-prepared CAD model of a link that is available on Onshape – an online CAD system platform. Anyone can access, edit, export, and download it. For more details and access, see [16], where the example Bennett mechanism files, ready for 3D printing, were also uploaded. The method get_design returns a tuple of Denavit-Hartenberg (DH) parameters, with proper automatic placement [17] of coordinate frames following the standard DH convention, and a set of Connection Point-parameter pairs for every link. The length parameters of the Bennett linkage from Figure 1 (right), scaled by 200 [-], result in the output of Table 1.
@@
As mentioned in the introduction, the 4 poses interpolation method from [7] is implemented in the package to synthesize 6R linkages with 1 DoF. The technique is based on computing a cubic motion polynomial interpolating the given poses. However, as explained in the paper, such a polynomial need not always to exist. Additionally, the authors impose restrictions on the given poses to avoid planar or spherical mechanisms and prismatic joints. We refer to [7] for potential pitfalls in case the interpolation in the Rational Linkages package does not yield satisfactory results.
@@ 4 Conclusion
This paper briefly presents a new Python-based package that implements various algorithms to deal with rational single-loop linkage synthesis and design, with the intention of simplifying their prototyping. Ready-to-run Python scripts and extended information on the presented examples, including results that were too long to present here, may be found in the documentation page [13] specially created for this paper.
Future development will be focused on collision-free realizations of these linkages, which is a topic that has not been much investigated. The algorithm by Li, Nawratil, et al. [1] which almost always finds a collision-free line model of any given 6R mechanism is also planned to be implemented. Furthermore, curved links can be applied, as in [18, 19], which can greatly expand the design possibilities, or the connection to a multibody dynamics system, as Exudyn [20], may be developed. This will bring closed-loop linkages much closer to engineering applications, serving, for example, as cheap single-purpose devices in robotic manipulation.

### v2

# references.bib

@@ 1 Introduction
Single-loop N-bar linkages are closed kinematic chains that connect a number of rigid bars via revolute or prismatic joints in one loop. Over-constrained mechanisms are a special class of such linkages. In this class, the Grübler-Kutzbach-Chebyshev formula, computing the degrees-of-freedom (DoF) of the mechanism, fails and the causes of this failure are special geometric properties of the design parameters of the mechanism. Because of the low number of DoFs, single-loop linkages can be synthesized as simple single-purpose devices for specific tasks, such as pick-and-place operations. This can be advantageous compared to serial robots, because they can perform complicated motions following space-curves of higher degree with a low number of active joints. Up to now, a major problem for the industrial application of these mechanisms is self-intersections [Li2020], as discussed later.
Custom mechanisms designed for specific tasks are a wide research topic, with many studies conducted in fields such as robot manipulation [Yihun2014, Dogra2022, Huczala2022] or locomotion [Xi2018, Zhang2019], whereby the design parameters are mostly obtained by numerical optimization. The tools presented in this paper, however, rely on algebraic methods. The problem of designing a spatial linkage analytically can be divided into the following steps: Pose interpolation, synthesis of the kinematic structure, and design of its physical realization, which also involves a collision analysis and CAD modeling to prepare 3D printing.
@@
The prime tasks in designing a custom mechanism or manipulator is the determination of the poses, which should be achieved by the end-effector tool of the device. The presented package implements for this purpose a method introduced by Hegedüs et al. [Hegeds2015], which interpolates four poses using a rational curve in the special Euclidean displacement group SE(3). When a rational motion curve on SE(3) is obtained, rational motion factorization introduced by Hegedüs, Schicho, and Schröcker [Hegeds2012, Hegeds2013] is applied. The factors of the rational motion curve relate to 1-DoF revolute or prismatic joints, which can be connected to form a linkage capable of performing the synthesized end-effector motion.
@@ 2 Mathematical Background
A possibility to describe lines in 3D space using coordinates are the so called Plücker coordinates. These are six-dimensional vectors $\mathbf{l}=(g_{0},g_{1},g_{2},g_{3},g_{4},g_{5})$. The line’s direction is $\mathbf{g}=(g_{0},g_{1},g_{2})$, its moment vector $\overline{\mathbf{g}}=(g_{3},g_{4},g_{5})$ is obtained as $\overline{\mathbf{g}}=\mathbf{q}\times\mathbf{g}$ where $\mathbf{q}$ is an arbitrary point on the line. Plücker coordinates fulfill the Plücker condition $\mathbf{g}^{T}\cdot\overline{\mathbf{g}}=0$. More on Plücker coordinates can be found in Pottmann and Wallner [Pottmann2001, Chapter 2].
Study parameters or dual quaternions are used for the description of rigid body transformations either as 8-tuples $\mathbf{p}=(p_{0},p_{1},\ldots,p_{7})$ or by using two quaternions $\mathbf{a},\mathbf{b}$ as dual quaternion $\mathbf{p}=\mathbf{a}+\varepsilon\mathbf{b}$, where $\mathbf{a}=(p_{0},p_{1},p_{2},p_{3})$, $\mathbf{b}=(p_{4},p_{5},p_{6},p_{7})$ and $\varepsilon$ is the dual unit $\varepsilon^{2}=0$ (see Bottema and Roth [bottema1979, Chapter XIII]). Study parameters have to fulfill the quadratic equation $\mathbf{a}^{T}\cdot\mathbf{b}=0$, which is the equation of the so-called Study quadric. Points $\mathbf{q}$ and lines $\mathbf{l}$ can be embedded into the algebra of dual quaternions by $\mathbf{q}=(1,0,0,0,0,q_{x},q_{y},q_{z})$, where $(q_{x},q_{y},q_{z})$ are the coordinates of the point in the Euclidean space $E^{3}$ and $\mathbf{l}=(0,g_{0},g_{1},g_{2},0,g_{3},g_{4},g_{5})$ where $g_{i}$ are the Plücker coordinates of the line. Transformation $\mathbf{p}$ transforms a point or a line by
@@
Given a rigid body motion represented by the polynomial $C(t)$ over the dual quaternions, it is generically possible to synthesize a linkage that performs this motion by combining different factorizations of $C(t)$, c.f. Hegedüs, Schicho, and Schröcker [Hegeds2013].
We compute the factorizations of the motion polynomial with Python using the open source package biquaternion-py by Thimm [bq-link-docs]. This package provides a user-friendly implementation of a general dual quaternion algebra and polynomials over it for numerical and symbolical computations. It is ready for use, even by non-specialists but some attention needs to be paid to details of symbolic computations (providing extension fields over which polynomials factor if they cannot be inferred implicitly) and numerics (for example taking into account rounding errors in polynomial division). More about the general usage of this package and its intricacies can be found in the documentation of the biquaternion-py package [bq-link-docs].
@@ 3 Rational Mechanism Toolbox
While the previous section introduced methods and tools that are implemented but are already known, this section presents the main contribution of this work. It is the added functionality in terms of rapid prototyping of the linkages. Installation instructions can be found in the documentation hosted on the ReadTheDocs platform and are accessible at [rat-link-docs]. The source is available from Gitlab repository hosted by the University of Innsbruck [rat-link-gitlab], which also serves as the code maintenance server.
As an example to demonstrate the features, we will use a general Bennett mechanism that was synthesized by Brunnthaler et al. [brunnthaler2005new]. The motion curve of degree two of the given Bennett linkage has the equation specified by Study parameters
@@
c = RationalCurve(C(t)) $\triangleright$ create an object Rational Curve from the input equations
factors = FactorizationProvider.factorize_motion_curve(c) $\triangleright$ factorize curve
m = RationalMechanism(factors) $\triangleright$ create a mechanism object
plt = Plotter.plot$(m)$triangleright plot the mechanism
@@
In the toolbox, a collision check for the line model is implemented. It uses dual quaternion operations on lines to create parametric equations of moving lines, joints and links, during a full cycle of the motion. The lines move relatively between each other and two lines $\mathbf{l_{0}}$ and $\mathbf{l_{1}}$ represented by their Plücker coordinates intersect if the following equation is fulfilled [Pottmann2001, p.140]:
@@
In the package, the command RationalMechanism.get_design(scale) generates design parameters for a pre-prepared CAD model of a link that is available on Onshape – an online CAD system platform. Anyone can access, edit, export, and download it. For more details and access, see [bennett-zenodo], where the example Bennett mechanism files, ready for 3D printing, were also uploaded. The method get_design returns a tuple of Denavit-Hartenberg (DH) parameters, with proper automatic placement [Huczala2022iccma] of coordinate frames following the standard DH convention, and a set of Connection Point-parameter pairs for every link. The length parameters of the Bennett linkage from Figure 1 (right), scaled by 200 [-], result in the output of Table 1.
@@
As mentioned in the introduction, the 4 poses interpolation method from [Hegeds2015] is implemented in the package to synthesize 6R linkages with 1 DoF. The technique is based on computing a cubic motion polynomial interpolating the given poses. However, as explained in the paper, such a polynomial need not always to exist. Additionally, the authors impose restrictions on the given poses to avoid planar or spherical mechanisms and prismatic joints. We refer to [Hegeds2015] for potential pitfalls in case the interpolation in the Rational Linkages package does not yield satisfactory results.
@@ 4 Conclusion
This paper briefly presents a new Python-based package that implements various algorithms to deal with rational single-loop linkage synthesis and design, with the intention of simplifying their prototyping. Ready-to-run Python scripts and extended information on the presented examples, including results that were too long to present here, may be found in the documentation page [rat-link-docs] specially created for this paper.
Future development will be focused on collision-free realizations of these linkages, which is a topic that has not been much investigated. The algorithm by Li, Nawratil, et al. [Li2020] which almost always finds a collision-free line model of any given 6R mechanism is also planned to be implemented. Furthermore, curved links can be applied, as in [Kot2020, Mlotek2023], which can greatly expand the design possibilities, or the connection to a multibody dynamics system, as Exudyn [Gerstmayr2023], may be developed. This will bring closed-loop linkages much closer to engineering applications, serving, for example, as cheap single-purpose devices in robotic manipulation.